### v1

# UTG: Towards a Unified View of Snapshot and Event Based Models for Temporal Graphs

Abstract
Temporal graphs have gained increasing importance due to their ability to model dynamically evolving relationships. These graphs can be represented through either a stream of edge events or a sequence of graph snapshots. Until now, the development of machine learning methods for both types has occurred largely in isolation, resulting in limited experimental comparison and theoretical cross-pollination between the two. In this paper, we introduce Unified Temporal Graph (UTG), a framework that unifies snapshot-based and event-based machine learning models under a single umbrella, enabling models developed for one representation to be applied effectively to datasets of the other. We also propose a novel UTG training procedure to boost the performance of snapshot-based models in the streaming setting. We comprehensively evaluate both snapshot and event-based models across both types of temporal graphs on the temporal link prediction task. Our main findings are threefold: first, when combined with UTG training, snapshot-based models can perform competitively with event-based models such as TGN and GraphMixer even on event datasets. Second, snapshot-based models are at least an order of magnitude faster than most event-based models during inference. Third, while event-based methods such as NAT and DyGFormer outperforms snapshot-based methods on both types of temporal graphs, this is because they leverage joint neighborhood structural features thus emphasizing the potential to incorporate these features into snapshot-based models as well. These findings highlight the importance of comparing model architectures independent of the data format and suggest the potential of combining the efficiency of snapshot-based models with the performance of event-based models in the future.
1 Introduction
Recently, Graph Neural Networks (GNNs)[9, 24] and Graph Transformers[31, 17] have achieved remarkable success in various tasks for static graphs, such as link prediction, node classification, and graph classification [6]. These successes are driven by standardized empirical comparisons across model architectures [6] and theoretical insights into the expressive power of these models [29].
However, real-world networks such as financial transaction networks [22], social networks [14], and user-item interaction networks [10] are constantly evolving and rarely static. These evolving networks are often modeled by Temporal Graphs (TGs), where entities are represented by nodes and temporal relations are represented by timestamped edges between nodes. Temporal graphs are categorized into two types: Discrete-Time Dynamic Graphs (DTDGs) and Continuous-Time Dynamic Graphs (CTDGs) [8]. DTDGs are represented by an ordered sequence of graph snapshots, while CTDGs consist of timestamped edge streams. Both representations of temporal graphs are prevalent in real-world applications.
Until now, the development of ML methods for both types has occurred mostly independently, resulting in limited experimental comparison and theoretical cross-pollination between the two. We argue that the time granularity of the data collection process together with the requirements of the downstream task have created a gap between DTDG and CTDG in model development and evaluation.
Isolated Model Development. Despite the similarities between DTDGs and CTDGs, models for these graphs have been developed largely in isolation. Adopting the terminology of [11], models targeting DTDGs focus on learning from a sequence of graph snapshots (snapshot-based models[30, 2, 15] , while methods for CTDGs focus on learning from a stream of timestamped edge events (event-based models[18, 16, 12]. The disparate data representations of DTDGs and CTDGs have impeded comprehensive comparison across models developed for each category. Consequently, there are limited theoretical insights and empirical evaluations of the true potential of these models when compared together. In real-world applications, representing the data as CTDGs or DTDGs is often a design choice, and the ambiguity of the actual performance merits of both categories makes it challenging to select the optimal model in a practical setting.
Distinct Evaluation Settings. Another obstacle to comparing snapshot and event-based methods is their distinct evaluation settings. Snapshot-based methods have been primarily tested under the deployed setting [30, 15], where the test set information is strictly not available to the model, and training set information is used for prediction. In contrast, event-based models are designed for the streaming setting [18, 7], where streaming predictions allow the model to use recently observed information, enabling event-based models to update their node representations at test time.
In this work, we aim to bridge the gap between event-based and snapshot-based models by providing a unified framework to train and evaluate them to predict future events on any type of temporal graph. Our main contributions are as follows:
  • Unified framework: We propose Unified Temporal Graph (UTG), a framework that unifies snapshot-based and event-based temporal graph models under a single umbrella, enabling models developed for one representation to be applied effectively to datasets of the other.
  • Updating snapshot-based models: We propose a novel UTG training strategy to boost the performance of snapshot-based models in the streaming setting. This allows snapshot-based models to achieve competitive performance with event-based models such as TGN and GraphMixer on the tgbl-wiki and Reddit CTDG datasets.
  • Benchmarking: By leveraging the UTG framework, we conduct the first systematic comparison between snapshot and event-based models on both CTDG and DTDG datasets. While some event-based methods such as NAT and DyGFormer outperform snapshot-based methods on both CTDGs and DTDGs, we posit this is due to leveraging joint neighborhood structural features rather than a fundamental property of event-based methods. Additionally, snapshot-based methods are at least an order of magnitude faster than event-based methods while achieving competitive performance. This suggests several future directions, such as integrating joint neighborhood structural features in snapshot-based models and developing a universal method that combines accuracy and efficiency for both DTDGs and CTDGs.
Reproducibility: The datasets used in this work are public. The code for reproducing the experiments are available in the supplementary materials.
2 Related Work
Recently, many ML methods were developed for temporal graphs. The well-adopted categorization by Kazemi et al. [8] considers two types of temporal graphs: Discrete Time Dynamic Graphs (DTDGs) and Continuous Time Dynamic Graphs (CTDGs). More detailed discussions on related work can be found in Appendix F.
Discrete Time Dynamic Graphs. Early methods often represent temporal graphs as a sequence of graph snapshots while adapting common graph neural networks such as Graph Covolution Network (GCN) [9] used in static graphs for DTDGs. For example, EGCN [15] employs a Recurrent Neural Network (RNN) to evolve the parameters of a GCN over time. In comparison, GCLSTM [2] learns the graph structure via a GCN while capturing temporal dependencies with an LSTM network [5]. Pytorch Geometric-Temporal (PyG Temporal) [20] is a comprehensive framework that facilitate neural spatiotemporal signal processing which implements existing work such as EGCN and GCLSTM in an efficient manner. HTGN [30] utilizes hyperbolic geometry to better capture the complex and hierarchical nature of the evolving networks. Recently, You et al. [32] introduced a novel live-update setting where GNNs are always trained on the most recent observed snapshot after making predictions. In comparison, the streaming setting in this work allows the model to use observed snapshots for forward pass but no training are permitted on test set.
Continuous Time Dynamic Graphs. Event-based methods process temporal graphs as a stream of timestamped edges. DyRep [23] and JODIE [10] are two pioneering work on CTDGs. TGAT [28] is one of the first work for studying inductive representation learning on temporal graphs. Rossi et al. [18] introduce Temporal Graph Networks (TGNs), a generic inductive framework of Temporal Graph Networks, showing DyRep, JODIE and TGAT as its special cases. Methods such as CAWN [26] and NAT [12] both focuses on learning the joint neighborhood of the two nodes of interest in the link prediction task. CAWN focuses on learning from temporal random walks while NAT is a neighborhood-aware temporal network model that introduces a dictionary-type neighborhood representation for each node. TCL [25] and DyGFormer [33] applies transformer based architecture on CTDGs, inspired by the success of transformer based architectures on time series [27], images [4] and natural language processing [1].
Evaluations in Graph Machine Learning. Datasets spanning a multitude of domains could have distinct underlying mechanisms of link formation; thus, it is important to explore link prediction across diverse datasets [13]. The Open Graph Benchmark [6] provides a large collection of datasets and tasks for benchmarking on static graphs. Recently, Huang et al. [7] proposed the Temporal Graph Benchmark (TGB), a collection of large-scale and realistic datasets from distinct domains for both link and node level tasks on CTDGs. In this work, we experiment on the tgbl-wiki and tgbl-review dataset to compare both snapshot and event-based methods. Poursafaei et al. [16] suggested that the performance of temporal link prediction is often over-optimistic due to the easy negative samples used. Thus, they propose to use the historical negatives which are past edges absent in the current step for evaluation. In addition, a simple heuristic baseline called Edgebank is proposed as a strong contender with fast computational time. In this work, we also incorporate historical negative samples for our evaluation and adding EdgeBank in the comparison.
3 Preliminaries
Definition 1 (Continuous Time Dynamic Graphs)
A Continuous Time Dynamic Graph (CTDG) $\mathcal{G}$ is formulated as a collection of edges represented as tuples with source node, destination node, start time and end time;
$\mathcal{G}=\{(s_{0},d_{0},t_{0}^{s},t_{0}^{e}),(s_{1},d_{1},t_{1}^{s},t_{1}^{e}),\dots,(s_{k},d_{k},t_{k}^{s},t_{k}^{e})\}$
where, for edge $i\in[0,k]$, $s_{i}$ and $d_{i}$ denote source and destination respectively. The start times are ordered chronologically $t_{0}^{s}\leq t_{1}^{s}\leq...\leq t_{k}^{s}$, each start time is less than or equal to the corresponding end time $t_{i}^{s}\leq t_{i}^{e}$, hence for each timestamp $t$ we have $t\in[t_{0}^{s},t_{k}^{e}]$.
Without loss of generality, one can normalize the timestamps in $\mathcal{G}$ from $[t_{0},t_{k}]$ to $[0,1]$ by applying $t=\frac{t-t_{0}}{t_{k}}\forall t\in[t_{0},t_{k}]$. Real world temporal networks can be broadly classified into two inherent types based on the nature of their edges: spontaneous networks and relationship networks. Examples of spontaneous networks include transaction networks, retweet networks, Reddit networks, and other activity graphs. Here, the edges are spontaneous thus resulting in the start time and end time of an edge being the same, i.e. $t^{s}=t^{e}$. This formulation is inline with related studies in [16, 8, 7, 21]. For relationship networks such as friendship networks, contact networks, and collaboration networks, the edges often persist over a period of time resulting in $t^{s}\neq t^{e}$.
Definition 2 (Discrete Time Dynamic Graphs)
A Discrete Time Dynamic Graph (DTDG) $\mathbf{G}$ is a sequence of graph snapshots sampled at regularly-spaced time intervals [8]:
$\mathbf{G}=\{\mathbf{G}_{0},\mathbf{G}_{1},\dots,\mathbf{G}_{T}\}$
$\mathbf{G}_{t}=\{\mathbf{V}_{t},\mathbf{E}_{t}\}$ is the graph at snapshot $t\in[0,T]$, where $\mathbf{V}_{t}$, $\mathbf{E}_{t}$ are the set of nodes and edges in $\mathbf{G}_{t}$, respectively.
4 UTG Framework
Figure 1: Illustration of the UTG framework. The input graph is processed by the UTG input mapper to generate the appropriate input data format for event-based or snapshot-based models. The model predictions are then processed by the UTG output mapper for the respective prediction task.
In this section, we present the Unified Temporal Graph (UTG) framework which aims to unify snapshot-based and event-based temporal graph models under the same framework, enabling temporal graph models to be applied to both CTDGs and DTDGs. UTG has two key components: input mapper and output mapper. Input mapping converts the input temporal graph into the appropriate representation needed for a given method, i.e. snapshots or events. Output mapper transforms the prediction of the model to the required time granularity of the task. Figure 1 shows the workflow of UTG framework. UTG enables any temporal graph learning methods to be applied to any input temporal graph via UTG input mapper and UTG output mapper.
4.1 UTG Input Mapper
Both snapshot and event-based TG methods require specific input data format. For snapshot-based models, discretizing CTDG data into a sequence of snapshots is required (Section 4.1.1). For event-based models, DTDG snapshots need to be converted into batches of events (Section 4.1.2).
4.1.1 Converting CTDG to Snapshots
Here, we formulate the discretization process which converts a continuous-time dynamic graph into a sequence of graph snapshots for snapshot-based models.
Definition 3 (Discretization Partition)
Let 00 and $1$ be the normalized start and end time of a temporal graph $\mathcal{G}$. A discretization partition $\mathbf{P}$ of the interval $[0,1]$ is a collection of intervals:
$\mathbf{P}=\{[\tau_{0},\tau_{1}],[\tau_{1},\tau_{2}],\dots,[\tau_{k-1},\tau_{k}]\}$
such that $10=\tau_{0}<\tau_{1}<\dots<\tau_{k}=$ 1 and where $k\in\mathbb{N}$.
Hence, a discretization partition $\mathbf{P}$ defines a finite collection of non-overlapping intervals and its norm is defined as:
$||\mathbf{P}||=\max\{|\tau_{1}-\tau_{0}|,|\tau_{2}-\tau_{1}|,\dots,|\tau_{k}-\tau_{k-1}|\}$
The norm $||\mathbf{P}||$ can also be interpreted as the max duration of a snapshot in the temporal graph $\mathcal{G}$. The cardinality of $\mathbf{P}$ is denoted by $|\mathbf{P}|$.
Definition 4 (Regular Discretization Partition)
A given discretization partition $\mathbf{P}$ is regular if and only if:
$\forall[\tau_{i},\tau_{j}]\in\mathbf{P},\;|\tau_{j}-\tau_{i}|=||\mathbf{P}||=\frac{|\tau_{k}-\tau_{0}|}{|\mathbf{P}|}$
In this case, all intervals have the same duration equal to $||\mathbf{P}||$.
Definition 5 (Induced Graph Snapshots)
Given a Continuous Time Dynamic Graph $\mathcal{G}$ and a Regular Discretization Partition $\mathbf{P}$, the Induced Graph Snapshots $\mathbf{G}$ are formulated as:
$\mathbf{G}=\{\mathbf{G}_{\tau_{0}}^{\tau_{1}},\mathbf{G}_{\tau_{1}}^{\tau_{2}},\dots,\mathbf{G}_{\tau_{k-1}}^{\tau_{k}}\}$
where $\mathbf{G}_{\tau_{i}}^{\tau_{j}}$ is defined as the aggregated graph snapshot containing all edges that have a start time $t_{s}\leq\tau_{j}$ and an end time $t_{e}\geq\tau_{i}$, i.e. edges that are present solely within the $[{\tau_{i}},{\tau_{j}}]$ interval.
Note that for spontaneous networks, each edge exist at a specific time point $t_{s}=t_{e}$ thus only belonging to a single interval/snapshot. For relationship networks however, it is possible for an edge to belong to multiple intervals depending on its duration.
Definition 6 (Discretization Level)
Given a regular discretization partition $\mathbf{P}$ and the timestamps in a temporal graph $\mathcal{G}$ normalized to $[0,1]$, the discretization level $\Delta$ of $\mathbf{P}$ is computed as :
$\Delta=\frac{1}{|\mathbf{P}|}$
where $|\mathbf{P}|$ is the cardinality of the partition or the number of intervals.
Note that $\Delta\in[0,1]$. When $|\mathbf{P}|=1$ then $\Delta=1$ which means the temporal graph is collapsed into a single graph snapshot (i.e. a static graph). On the other extreme, we have $\lim_{|\mathbf{P}|\to\infty}\Delta=0$, preserving the continuous nature of the continuous time dynamic graph $\mathcal{G}$.
Definition 7 (Time Gap)
Given a Continuous Time Dynamic Graph $\mathcal{G}$ and a Regular Discretization Partition $\mathbf{P}$, a Time Gap occurs when there exist one or more snapshots in the Induced Graph Snapshots $\mathbf{G}$ with an empty edge set.
In this work, we choose the number of intervals in discretization by selecting the finest time granularity which would not induce a time gap. This ensures that there are no empty snapshots in the induced graph snapshots.
4.1.2 Converting DTDG to Events
While it may seem straightforward to convert DTDG to events — simply create one event with timestep $t$ for each edge in snapshot $G_{t}$ — some subtleties related to batch training and memory update of event-based models have to be considered to avoid data leakage. Event-based models often receive batches of events (or edges) with a fixed dimension as inputs [18, 33, 12]. In discrete-time dynamic graphs, all edges in a snapshot have the same timestamp and are assumed to arrive simultaneously. Therefore, using a fixed batch size can result in splitting the snapshot into multiple batches. Because models in the streaming setting [7] such as TGN [18] and NAT [12] update their representation of the temporal graph at the end of each batch, predicting a snapshot across multiple batches leads to data leakage: a portion of the edges from the snapshot is used to predict other (simultaneous) edges from the same snapshot. To avoid data leakage on DTDGs, we ensure that each snapshot is contained in a single batch for event-based models [^1] In case the resulting batch would not fit in memory, one can delay the memory update (and parameter updates during training) only after all edges from the current snapshot have been processed, something akin to gradient accumulation..
4.2 UTG Output Mapper
The output task on the temporal graph can be either discrete or continuous. Discrete tasks refer to predicting which edges will be present at a future snapshot (with an integer timestep). Continuous tasks refer to predicting which edges will be present for a given UNIX timestamp in the future. Snapshot-based models often omits the timestamp of the prediction as an input, implicitly assuming the prediction is for the next snapshot. Therefore, applying snapshot-based models for a continuous task requires 1). always updating the model with all the information available until the most recent observed snapshot (test-time update) and 2). mapping the discrete-time prediction to a continuous timestamp. We explain here how to map the prediction to a continuous space with zero-order hold.
Definition 8 (Zero-order Hold)
A discrete time signal $y[i]$, $i\in\mathbb{N}$, can be converted to a continuous time signal $y(t)$, $t\in\mathbb{R}$, by broadcasting the value y[i] as a constant in the interval $[\tau_{i},\tau_{j}]$:
$y(t)=y[i],\ \ \text{for all}\quad\tau_{i}\leq t\leq\tau_{j}$
where $[\tau_{i},\tau_{j}]$ specifies the duration of the discrete signal.
By applying zero-order hold for snapshot-based models, the predictions can now be broadcasted for a period of time (specifically for the duration of a given snapshot $[\tau_{i},\tau_{j}]$). Therefore, it is now possible to utilize snapshot-based models on continuous-time dynamic graphs.
5 Improving Snapshot-based Models
Here, we discuss the changes to the snapshot-based methods in the UTG framework.
5.1 Streaming Setting
Figure 2: UTG training and evaluation workflow. UTG training enables snapshot-shot based models to incorporate information from observed snapshots at inference time.
Event-based models often evaluate with the streaming setting [18, 33, 26, 12]. In this setting, information from the previously observed batches of events (or graph snapshot) can be used to update the model however no information from the test set is used to train the model.
We also adopt the streaming setting which closely resembles real-world settings where after the model is trained, it is required to incorporate newly observed information into its predictions. Figure 2 shows the evaluation pipeline used for snapshot-based models in UTG. After a snapshot is observed, it can be used to update the node representation of the snapshot-based models for the prediction of the next snapshot (only forward pass for inference). More discussion on evaluation settings are shown in Appendix E.
5.2 UTG Snapshot Training
Figure 2 illustrates the UTG training procedure for snapshot-based models. In standard training, e.g. with Pytorch Geometric Temporal [19], the model uses snapshots up to $\mathbf{G}_{t}$ to predict edges for the same time $t$, leading to information leakage as the model predicts edges that were part of its input. Moreover, the gradients from all training snapshots are accumulated before performing backpropagation to update the model weights. In contrast, UTG training introduces two key differences: (1) only snapshots up to $\mathbf{G}_{t-1}$ are used to predict the edges for time $t$, and (2) it performs backpropagation at each snapshot rather than accumulating gradients for all training snapshots. In Section 6.4, we conduct an ablation study on the GCLSTM model to demonstrate the benefits of UTG training.
6 Experiments
In this section, we benchmark both snapshot-based and event-based methods across both CTDG and DTDG datasets under the UTG framework.
Datasets. In this work, we consider five discrete-time dynamic graph datasets and three continuous-time dynamic graph datasets. tgbl-wiki and tgbl-review are datasets from TGB [7] while the rest are found in Poursafaei et al. [16]. The dataset statistics are shown in Table 1. The time granularity or discretization level of each DTDG dataset is selected as the finest time granularities where there are no time gaps (see Section 4.1.1). The surprise index is defined as $surprise=\frac{|E_{\text{test}}\setminus E_{\text{train}}|}{E_{\text{test}}}$ [16] which measures the proportion of unseen edges in the test set when compared to the training set.
Table 1: Dataset statistics.
Dataset # Nodes # Edges # Unique Edges Surprise Time Granularity # Snapshots
DTDG UCI 1,899 26,628 20,296 0.535 Weekly 29
Enron 184 10,472 3,125 0.253 Monthly 45
Contact 694 463,558 79,531 0.098 Hourly 673
Social Evo. 74 87,479 4,486 0.005 Daily 244
MOOC 7,144 236,808 178,443 0.718 Daily 31
CTDG tgbl-wiki 9,227 157,474 18,257 0.108 UNIX timestamp 745 (Hourly)
tgbl-review 352,637 4,873,540 4,730,223 0.987 UNIX timestamp 237 (Monthly)
Reddit 10,984 672,447 78,516 0.069 UNIX timestamp 745 (Hourly)
Evaluation Setting. For each dataset, we generate a fixed set of negative samples for each positive edge consisting of 50% historical negatives and 50% random negative, same as in TGB [7]. For DTDG datasets, we generate 1000 negative samples per positive edge. For TGB datasets, we use the same set of negatives provided in TGB and for Reddit, we generate 1000 negatives similar to before. For graphs with less than 1k nodes, we generate negative samples equal to number of nodes. We follow the streaming setting where the models are allowed to update their representation at test time while gradient updates are not permitted. For DTDG datasets, we select the best results from learning rate $0.001$ or $0.0002$. For CTDG datasets, we report the results from TGB [7] where available or by learning rate $0.0002$.
Compared Methods. We compare four event-based methods including TGN [18], DyGFormer [33], NAT [12] and GraphMixer [3]. We also include Edgebank [16], a scalable and non-parameteric heuristics. In addition, we compare three existing snapshot-based methods including HTGN [30], GCLSTM [2] and EGCNo [15]. Lastly, we adapt a common 2-layer (static) GCN [9] under the UTG framework to demonstrate the flexibility of UTG (without a recurrent module). If a method runs out of memory on a NVIDIA A100 GPU (40GB memory), it is reported as out of memory (OOM). If a method runs for more than 5 days, it is reported as out of time (OOT)
6.1 Comparing Event-Based with Snapshot-Based Models
With the UTG framework, we can now compare snapshot-based methods and event-based methods on any temporal graph dataset. This comparison allows us to focus on analyzing the strength and weaknesses of the model design, independent of the data format.
Table 2: Test MRR comparison for snapshot and event-based methods on DTDG datasets, results reported from 5 runs. Top three models are coloured by First, Second, Third.
{tabu}
l | l | l l l l l
Method UCI Enron Contacts Social Evo. MOOC
event TGN [18] 0.091 $\pm$ 0.002 0.191 $\pm$ 0.027 0.153 $\pm$ 0.007 0.283 $\pm$ 0.009 0.174 $\pm$ 0.009
DyGFormer [33] 0.334 $\pm$ 0.024 0.331 $\pm$ 0.010 0.283 $\pm$ 0.006 0.366 $\pm$ 0.004 OOM
NAT [12] 0.356 $\pm$ 0.048 0.276 $\pm$ 0.014 0.245 $\pm$ 0.015 0.258 $\pm$ 0.036 0.283 $\pm$ 0.058
GraphMixer [3] 0.105 $\pm$ 0.008 0.296 $\pm$ 0.019 0.055 $\pm$ 0.003 0.157 $\pm$ 0.005 OOM
$\text{EdgeBank}_{\infty}$ [16] 0.055 0.115 0.016 0.049 0.040
$\text{EdgeBank}_{\text{tw}}$ [16] 0.165 0.157 0.050 0.070 0.070
snapshot HTGN (UTG) [30] 0.093 $\pm$ 0.012 0.267 $\pm$ 0.007 0.165 $\pm$ 0.001 0.228 $\pm$ 0.003 0.093 $\pm$ 0.005
GCLSTM (UTG) [2] 0.093 $\pm$ 0.006 0.170 $\pm$ 0.008 0.128 $\pm$ 0.004 0.286 $\pm$ 0.003 0.143 $\pm$ 0.006
EGCNo (UTG) [15] 0.121 $\pm$ 0.010 0.233 $\pm$ 0.008 0.192 $\pm$ 0.001 0.253 $\pm$ 0.006 0.126 $\pm$ 0.009
GCN (UTG) [9] 0.068 $\pm$ 0.009 0.164 $\pm$ 0.011 0.104 $\pm$ 0.002 0.289 $\pm$ 0.008 0.084 $\pm$ 0.010
DTDG Results. Table 6.1 shows the performance of all methods on the DTDG datasets. Surprisingly, we find that event-based methods achieve state-of-the-art performance on the DTDG datasets, particularly with DyGFormer and NAT consistently outperforming other methods.
With the improvements from UTG, snapshot-based models can obtain competitive performance on datasets such as Enron and Social Evo. Interestingly, even the simple GCN with UTG training can achieve second place performance on the Social Evo. dataset. Note that this dataset has the lowest surprise out of all datasets meaning the majority of test set edges have been observed during training, possibly explaining the strong performance of GCN in this case. Lastly, on the MOOC dataset which has the largest number of nodes out of all DTDG datasets, both DyGformer and GraphMixer ran out of memory (OOM) showing their difficulty in scaling with the number of nodes in a snapshot.
Table 3: Test MRR comparison for snapshot and event-based methods on CTDG datasets, results reported from 5 runs. Top three models are coloured by First, Second, Third.
{tabu}
l | l | l l l
Method tgbl-wiki tgbl-review Reddit
event TGN [18] 0.396 $\pm$ 0.060 0.349 $\pm$ 0.020 0.499 $\pm$ 0.011
DyGFormer [33] 0.798 $\pm$ 0.004 0.224 $\pm$ 0.015 OOT
NAT [12] 0.749 $\pm$ 0.010 0.341 $\pm$ 0.020 0.693 $\pm$ 0.015
GraphMixer [3] 0.118 $\pm$ 0.002 0.521 $\pm$ 0.015 0.136 $\pm$ 0.078
$\text{EdgeBank}_{\infty}$ [16] 0.495 0.025 0.485
$\text{EdgeBank}_{\text{tw}}$ [16] 0.571 0.023 0.589
snapshot HTGN (UTG) [30] 0.464 $\pm$ 0.005 0.104 $\pm$ 0.002 0.533 $\pm$ 0.007
GCLSTM (UTG) [2] 0.374 $\pm$ 0.010 0.095 $\pm$ 0.002 0.467 $\pm$ 0.004
EGCNo (UTG) [15] 0.398 $\pm$ 0.007 0.195 $\pm$ 0.001 0.321 $\pm$ 0.009
GCN (UTG) [9] 0.336 $\pm$ 0.009 0.186 $\pm$ 0.002 0.242 $\pm$ 0.005
CTDG Results. Table 6.1 shows the performance of all methods on the CTDG datasets. Similar to DTDG datasets, DyGformer and NAT retains competitive performance here. On the tgbl-wiki and Reddit dataset, HTGN, a snapshot-based model is able to outperform widely-used TGN and GraphMixer model. This shows that learning from the discretized snapshots can be effective even on CTDG datasets. However, snapshot-based models have lower performance on the tgbl-review dataset where the surprise index is high. This shows that the inductive reasoning capability on snapshot-based models should be further improved to generalize to unseen edges.
6.2 Computational Time Comparison
(a) Test time (seconds) for Social Evo. dataset.
(b) Test time (seconds) for tgbl-review dataset.
Figure 3: Snapshot-based models are at least an order of magnitude faster than event-based models for inference.
Figure 2(a) and Figure 2(b) shows the test inference time comparison for all methods on the Social Evo. and the tgbl-review dataset respectively. The test time for each dataset is reported in Appendix D. Overall, we observe that snapshot-based methods are at least an order of magnitude faster than most event-based methods. In comparison, high performing model such as DyGformer has significantly higher computational time thus limiting its scalability to large datasets. One important future direction is to combine the predictive power of event-based models such as NAT and DyGFormer with that of the efficiency of the snapshot-based model for improved scalability.
6.3 Discussion
Event-based methods such as NAT and DyGFormer tend to perform best on both CTDG and DTDG, potentially leading to the premature conclusion that event-based modeling is the preferred paradigm and that snapshot-based models should be avoided. However, the superior performance of NAT and DyGFormer could be primarily due to their ability to leverage joint neighborhood structural features, specifically the common neighbors between the source and destination nodes of a link [12, 33]. This approach has been shown to be fundamental for achieving competitive link prediction on static graphs [34, 35]. In contrast, none of the existing snapshot-based methods incorporate joint neighborhood structural features. Therefore, the performance difference could be mainly attributed to this factor rather than an intrinsic difference between event-based and snapshot-based models. This is confirmed by the fact that event-based models such as TGN and GraphMixer, which omits these features, have no clear performance advantage over snapshot-based methods. Moreover, snapshot-based methods are more computationally efficient and might be preferred when efficiency is important. These considerations suggest that both event-based and snapshot-based methods have their own merit. We believe that combining the strengths of both approaches will be an important direction for future research.
6.4 UTG Training Ablation Study
Table 4: GCLSTM ablation with UTG training when compared with Pytorch Geometric-Temporal training, results averaged across five runs best results are bolded.
Dataset GCLSTM (UTG) GCLSTM (original)
UCI 0.093 $\pm$ 0.006 0.047 $\pm$ 0.006
Enron 0.170 $\pm$ 0.006 0.131 $\pm$ 0.003
Contacts 0.128 $\pm$ 0.004 0.101 $\pm$ 0.031
Social Evo. 0.286 $\pm$ 0.003 0.287 $\pm$ 0.009
MOOC 0.143 $\pm$ 0.006 0.076 $\pm$ 0.003
Table 4 compares the performance of GCLSTM using UTG training with that of Pytorch Geometric Temporal training on the DTDG datasets (Section 5.2). UTG training significantly enhances GCLSTM performance across four datasets and performs identically on the Social Evo dataset. This demonstrates the benefits of predicting the next snapshot and backpropagating after every snapshot.
7 Conclusion
In this work, we introduce the UTG framework, unifying both snapshot-based and event-based temporal graph models under a single umbrella. With the UTG input mapper and UTG output mapper, temporal graph models developed for one representation can be applied effectively to datasets of the other. To compare both types of methods in the streaming setting for evaluation, we propose the UTG training to boost the performance of snapshot-based models. Extensive experiments on five DTDG datasets and three CTDG datasets are conducted to comprehensively compare snapshot and event-based methods. We find that top performing models on both types of datasets leverage joint neighborhood structural features such as the number of common neighbors between the source and destination node of a link. In addition, snapshot-based models can achieve competitive performance to event-based model such as TGN and GraphMixer while being an order of magnitude faster in inference time. Thus, an important future direction is to combine the strength of both types of methods to achieve high performing and scalable temporal graph learning methods.
Appendix A Broader Impact
We expect our proposed UTG framework to impact the field of temporal graph learning in a positive way by allowing unified comparison between previously disjoint lines of research. The empirical findings from this paper suggest the potential to combine the advantages from both snapshot and event-based methods to design effective and scalable temporal graph methods in the future. The UTG framework can also be applied to any existing and future snapshot or event-based methods to enable them to evaluate on both types of temporal graph representations.
Appendix B Potential Negative Impact
Our proposed UTG is a framework which unifies different types of methods from temporal graph learning thus there are no direct negative impacts. However, temporal graph learning can be widely applied in applications such as fraud detection, traffic forecasting, recommendation systems and more. Without proper understanding of the behavior of TG methods, it is possible that the model predictions might be incorrect or misleading thus causing negative impact.
Appendix C Limitations
In this work, we focus on the streaming setting, which is the most commonly used setting for evaluation of event-based methods. However, other settings such as the deployed setting and the live-update setting as discussed in Appendix E are not considered due to computational constraints. Future work can utilize the UTG framework under the aforementioned settings.
Appendix D Computational Time
Table 5: Test inference time comparison for snapshot and event based methods on DTDG datasets, we report the average result from 5 runs. Top three models are coloured by First, Second, Third.
{tabu}
l | l | l l l l l
Method UCI Enron Contacts Social Evo. MOOC
event TGN [18] 1.07 1.71 137.57 24.04 50.04
DyGFormer [33] 155.58 57.72 15423.99 349.22 OOM
NAT [12] 3.82 8.39 596.22 148.43 299.00
GraphMixer [3] 32.88 13.85 3542.88 132.39 OOM
$\text{EdgeBank}_{\infty}$ [16] 0.52 0.24 45.33 2.07 5.17
$\text{EdgeBank}_{\text{tw}}$ [16] 0.52 0.25 50.77 2.45 6.12
snapshot HTGN (UTG) [30] 0.61 0.87 76.64 14.59 28.64
GCLSTM (UTG) [2] 0.35 0.46 40.83 9.27 19.78
EGCNo (UTG) [15] 0.43 0.45 40.62 7.35 15.49
GCN (UTG) [9] 0.50 0.31 56.88 6.40 13.30
Table 6: Test inference time comparison for snapshot and event based methods on CTDG datasets, results reported from 5 runs. Top three models are coloured by First, Second, Third.
{tabu}
l | l | l l l l
Method tgbl-wiki tgbl-review Reddit
event TGN [18] 39.24 1137.69 286.79
DyGFormer [33] 7196.52 26477.51 OOT
NAT [12] 340.51 8925.21 1159.19
GraphMixer [3] 1655.44 4167.63 7166.24
$\text{EdgeBank}_{\infty}$ [16] 20.06 140.25 26.91
$\text{EdgeBank}_{\text{tw}}$ [16] 20.67 143.49 27.08
snapshot HTGN (UTG) [30] 28.96 718.17 117.05
GCLSTM (UTG) [2] 20.54 436.30 82.88
EGCNo (UTG) [15] 20.15 433.23 84.85
GCN (UTG) [9] 18.25 384.51 78.78
Table D shows the inference time for all methods on DTDG datasets. Table D shows the inference time for all methods on CTDG datasets. OOM means out of memory and OOT means out of time. We observe that snapshot-based models are at least one order of magnitude faster than event-based models such as NAT, DyGFormer and GraphMixer. In addition, the best performing model on most datasets, DyGFormer, is also consistently the slowest method.
Appendix E Evaluation Settings
Figure 4: Different setting for evaluation of future link prediction include between deployed, streaming and live-update setting. UTG framework is designed for the streaming setting.
Deployed setting. The deployed setting is often used as the evaluation setting for snapshot-based methods [30, 15]. In this setting, no information from the test set is passed to the model, and the node embeddings from the last training snapshot are used for predictions in all test snapshots.
Streaming setting. event-based models often evaluate with the streaming setting [18, 33, 26, 12]. In this setting, information from the previously observed batches of events can be used to update the model however no information from the test set can be used to train the model.
Live-update Setting. You et al. [32] proposed the live-update setting where the model weights are constantly updated to newly observed snapshots while predicting the next snapshot. To predict links in $\mathbf{G}_{t+1}$, first the observed snapshot $\mathbf{G}_{t-1}$ are split into a training set and a validation set. The model is trained on $\mathbf{G}_{t-1}^{train}$ while using $\mathbf{G}_{t-1}^{val}$ for early stopping. Lastly, the trained model receives $\mathbf{G}_{t}$ and predicts for $\mathbf{G}_{t+1}$.
Figure 4 illustrates the difference between these three settings. In this work, we focus on the streaming setting as it closely resembles the common use case where even after a model is trained, it is expected to incorporate new information from the data stream for accurate predictions.
Appendix F Temporal Graph Learning Methods
F.1 Snapshot-based Methods
ROLAND [32]: You et al. [32] propose ROLAND as a framework for repurposing static GNNs to DTDGs. In ROLAND, a hierarchy of node states is constructed by different GNN layers that are recurrently updated over time. You et al. [32] introduce a novel evaluation setting, live-update, where GNNs are updated based on the recently observed events after making prediction. ROLAND provides the opportunity of adapting efficient static GNN designs for DTDGs.
HTGN [30]: HTGN utilizes hyperbolic geometry to better capture the complex and hierarchical nature of the evolving networks. It combines hyperbolic GNNs and hyperbolic gated recurrent network to jointly model temporal dynamics and hierarchical structures. HTGN consists of two main modules: hyperbolic temporal contextual self-attention that focuses on the historical state attention, and hyperbolic temporal consistency that provides stability and generalization.
GCLSTM [2]: GCLSTM learns the node structure s through a GCN, while the temporal features are captured with an LSTM. The models is proposed especially for dynamic link prediction task.
EGCN [15]: EvolveGCN focuses on dynamic graph representation learning for temporal networks where the set of nodes changes frequently. Basically, EvolveGCN exploits an RNN to adapt the parameters of a GCN over time. By focusing on adapting the model instead of the node embeddings, EvolveGCN is not restricted to the node set and can learn representation for the nodes without historical information.
PyG-Temporal [20]: PyTorch Geometric Temporal is a comprehensive framework that facilitate neural spatiotemporal signal processing. It offers a cohesive platform that simplifies temporal geometric deep learning. PyG Temporal is built upon the foundations of existing PyTorch libraries, and it includes streamlined neural network layer definitions, efficient temporal snapshot generators for batching, and a variety of benchmark datasets.
F.2 Event-based Methods
TGAT [28]: It is one of the pioneering work on inductive representation learning on temporal graphs. Xu et al. [28] argue that learning on dynamic graphs requires methods that handle new nodes and capture temporal patterns, with node embeddings representing both static features and evolving structures. The temporal graph attention (TGAT) layer is proposed to efficiently aggregate temporal-topological neighborhood features and learn time-feature interactions using self-attention. Stacking TGAT layers enables the model to treat node embeddings as functions of time, allowing inductive inference for new and existing nodes.
TGN [18]: Rossi et al. [18] introduce Temporal Graph Networks (TGNs), a versatile and efficient framework for dynamic graphs, represented as stream of timestamped events. TGNs leverage a combination of memory modules and graph-based operators to improve computational efficiency. Essentially, TGN is a framework that subsumes several previous models as specific instances.
CAWN [26]: To learn the underlying laws, such as triadic closure, that governs real-world network, Wang et al. [26] propose Causal Anonymous Walks (CAWs) for representing temporal networks inductively. CAWs, derived from temporal random walks, act as automatic retrievals of temporal network motifs, avoiding the need for their manual selection and counting. Wang et al. [26] alswo propose an anonymization strategy which replaces node identities with hitting counts from sampled walks, maintaining inductiveness and motif correlation. CAWN is a neural network model proposed to encode CAWs, paired with a CAW sampling strategy that ensures constant memory and time costs for online training and inference.
TCL [25]: TCL effectively learns dynamic node representations by capturing both temporal and topological information. It features three main components: a graph-topology-aware transformer adapted from the vanilla Transformer, a two-stream encoder that independently extracts temporal neighborhood representations of interacting nodes and models their interdependencies using a co-attentional transformer, and an optimization strategy inspired by contrastive learning. This strategy maximizes mutual information between predictive representations of future interaction nodes, enhancing robustness to noise.
NAT [12]: In modeling temporal networks, the neighborhood of nodes provides essential structural information for interaction prediction. It is often challenging to extract this information efficiently. Luo and Li [12] propose the Neighborhood-Aware Temporal (NAT) network model that introduces a dictionary-type neighborhood representation for each node. NAT records a down-sampled set of neighboring nodes as keys, enabling fast construction of structural features for joint neighborhoods. A specialized data structure called N-cache is designed to facilitate parallel access and updates on GPUs.
EdgeBank [16]: Poursafaei et al. [16] introduce stricter evaluation procedures for dynamic graph link prediction, reflecting real-world scenarios. They leverage the recurring patterns of edges to proposes EdgeBank that is a baseline relying on memorization with surprisingly good performance.
TGB [7]: This is a curated collection of comprehensive and varied benchmark datasets tailored for rigorous evaluation of machine learning models on temporal graphs. These datasets are extensive and diverse and the benchmark includes tasks for both node-level and edge-level prediction, with evaluation protocols designed around practical use cases. TGB provides an automated machine learning pipeline for reproducible and accessible research, encompassing data loading, experiment setup, and performance evaluation.
DyGFormer [33]: Yu et al. [33] introduces a transformer-based architecture for dynamic graph learning. DyGFormer focuses on learning from nodes historical first-hop interactions and employs a neighbor co-occurrence encoding scheme to capture correlations between source and destination nodes through their historical sequences. Yu et al. [33] also propose a patching technique that divides each sequence into patches for the transformer, enabling effective utilization of longer histories. Yu et al. [33] also present DyGLib that is a library for standardizing training pipelines, extensible coding interfaces, and thorough evaluation protocols to ensure reproducible dynamic graph learning research.
Appendix G Computing Resources
For our experiments, we utilized one of the following GPUs. The first option was NVIDIA A100 GPUs (40GB memory) paired with 4 CPU nodes. These nodes featured CPUs such as the AMD Rome 7532 @ 2.40 GHz with 256MB cache L3, AMD Rome 7502 @2.50 GHz with 128MB cache L3, or AMD Milan 7413 @ 2.65 GHz with 128MB cache L3, each equipped with 100GB memory. The second option was using NVIDIA V100SXM2 GPUs (16GB memory) alongside 4 CPU nodes, which housed Intel Gold 6148 Skylake CPUs @ 2.4 GHz, each with 100GB memory. Our last choice was to run experiments using NVIDIA P100 Pascal GPUs (12GB HBM2 memory) with 4 CPU nodes from Intel E5-2683 v4 Broadwell @ 2.1GHz with 100GB memory. Each experiment had a five-day time limit and was repeated five times, with results reported as averages and standard deviations. Notably, aside from methods adopted from the PyTorch Geometric library, several other models (assessed using their original source code or the DyGLib repository) encountered out-of-memory or out-of-time errors when applied to larger datasets.
Appendix H Model Configurations
For all methods and datasets, we employed the Adam optimizer with a two different learning rates namely $0.001$ and $0.0002$, and the configuration with the higher average performance was selected for reporting the results. Each experiment was repeated five times and the average and standard deviations were reported.
The train, validation, and test splits for tgbl-wiki and tgbl-review are provided by the TGB benchmark. For other datasets (namely, UCI, Enron, Contacts, Social Evo., MOOC, and Reddit), we used a chronological split of the data with $70\%$, $15\%$, and $15\%$ for the training, validation, and test set, respectively, which is inline with previous studies [16, 18, 28, 12]. We set the batch size equal to 64 for NAT, and for all other models (i.e., TGN, DyGFormer, GraphMixer, EdgeBank, HTGN, GCLSTM, EGCNo, and GCN) the batch size was 200. For the experiments on CTDGs, we set the number of epoch equal to 40 and implemented an early stopping approach with a patience of 20 epochs and tolerance of $10^{-5}$. For the experiments on DTDGs, the number of epochs was set to 200 with a similar early stopping approach. Dropout was set to 0.1. We set the number of attention heads equal to 2 for the models with an attention module, and node embedding size was fixed at 100. For TGN, the time embedding size was 100 and the memory dimension was specified as 172, with a message dimension of 100. For NAT, we set the bias=1e-5, and replacement probability=0.7. All other parameters were set according to the suggested values by Luo and Li [12]. The special hyperparameters of the DyGFormer and GraphMixer are set according to the recommendations presented by Yu et al. [33].